OpenStudy (anonymous):

If u = (1 + i ) and v = (1 − i ), what is uv

OpenStudy (anonymous):

use FOIL (1 + i)(1 - i) = 1 - i + i - i²) = 1 - i² = 1 - (-1) = 2